# MMTU: A Massive Multi-Task Table Understanding and Reasoning Benchmark

### v2

@@ Abstract
In this work, we introduce MMTU, a large-scale benchmark with over 30K questions across 25 real-world table tasks, designed to comprehensively evaluate models ability to understand, reason, and manipulate real tables at the expert-level. These tasks are drawn from decades’ worth of computer science research on tabular data, with a focus on complex table tasks faced by professional users. We show that MMTU require a combination of skills – including table understanding, reasoning, and coding – that remain challenging for today’s frontier models, where even frontier reasoning models like OpenAI o4-mini and DeepSeek R1 score only around 60%, suggesting significant room for improvement. We highlight key findings in our evaluation using MMTU and hope that this benchmark drives further advances in understanding and developing foundation models for structured data processing and analysis. Our code and data are available at https://github.com/MMTU-Benchmark/MMTU and https://huggingface.co/datasets/MMTU-benchmark/MMTU.
@@ 1 Introduction
Remarkable progress has been made in foundation models Bommasani et al. (2021); Brown et al. (2020); Guo et al. (2025); Touvron et al. (2023); Anil et al. (2023), partly thanks to an expanding array of large-scale benchmarking efforts. Prominent examples include benchmarks for general language understanding (e.g., GLUE Wang et al. (2018), Super-GLUE Wang et al. (2019), BIG-Bench Srivastava et al. (2022), MMLU Hendrycks et al. (2020), MMLU-pro Wang et al. (2024)), as well as benchmarks focused on coding and STEM reasoning (e.g., SWE-bench ben (i), GPQA-diamond Rein et al. (2024), AIME ben (a), LiveCodeBench ben (g)). These efforts have played a critical role in deepening our understanding and accelerating the progress of foundation models.
Tables and table-based use cases are central to many real-world applications, including spreadsheets exc (a); Goo , databases dat (b, c), and computational notebooks jup ; dat (d), which often require expert-level users such as data engineers, analysts, and database administrators to operate. While LLMs have shown great promise in working with tables exc (b); dat (a); gem , existing evaluations of table-tasks remain narrow in scope – primarily focusing on NL-to-SQL Li et al. (2024); Zheng et al. (2024); Yu et al. (2018) and Table-QA Yang et al. (2015); Cheng et al. (2021); Chen et al. (2019) – and fail to reflect the broader spectrum of real-world tasks that professional users face.
@@
In this work, we introduce MMTU, a large-scale and challenging benchmark comprising 30,647 questions across 25 real-world table tasks. It is designed to comprehensively evaluate models’ ability to understand, reason, and manipulate real tables at the expert-level. The tasks we collect in MMTU are drawn from decades of computer science research on tabular data, with a particular focus on complex tasks that professional users face, as illustrated in Figure 1.
Our evaluation shows that the complex and technical nature of MMTU demands a combination of capabilities – including table understanding, reasoning, and coding – that remain challenging for today’s frontier models. Even the top-performing models, such as OpenAI o4-mini and DeepSeek , achieve only 63.9% and 59.6% on MMTU, suggesting substantial room for improvement, and highlighting MMTU as a strong testbed for models aspiring toward general human-level intelligence, e.g., to meet or surpass the top 10% of skilled adults in diverse technical tasks like characterized in Morris et al. (2023).
@@
  • Reasoning-focused models, such as DeepSeek and OpenAI o4-mini, show a clear advantage over general-purpose chat models like GPT-4o and Llama. The top reasoning models outperform the top chat models by over 10 percentage points (Table 3), underscoring the complexity of the tasks (which often require coding in SQL/Pandas) and the importance of reasoning in MMTU.
@@
Table Transform table-transform-by-relationalization (TTBR) Relationalize a table using transformation Acc Li et al. (2023b); Barowy et al. (2015); Jin et al. (2017) 238
@@
table-transform-by-output-schema (TTBS) Synthesize transformation by output schema Acc Yang et al. (2021) 700
@@
table-transform-by-output-table (TTBT) Synthesize transformation by input/output tables Acc Bavishi et al. (2019); Wang et al. (2017); Tran et al. (2009); Zhang and Sun (2013) 87
@@
Table Matching Entity matching (EM) Match rows refer to the same semantic entity Acc Das et al. ; Mudgal et al. (2018); Zhao and He (2019); Li et al. (2020); Papadakis et al. (2021) 4380
@@
Schema matching (SM) Match columns refer to the same concept F1 Koutras et al. (2021); Zhang et al. (2024); Bellahsene et al. (2011); Madhavan et al. (2001) 723
@@
Head value matching (HVM) Match column-headers with cell-values Acc Li et al. (2023c); Rahm and Bernstein (2001) 952
@@
Data Cleaning data-imputation (DI) Predict missing values in tables Acc Cafarella et al. (2008a); Eggert et al. (2023); Li et al. (2023c) 2000
@@
error-detection (ED) Detect erroneous cells in tables F1 Chu et al. (2016); Wang and He (2019); Mahdavi and Abedjan (2020); Chen et al. (2025) 1987
@@
list-to-table (L2T) Split lists of undelimited values into table Acc Chu et al. (2015); Cafarella et al. (2008b); Elmeleegy et al. (2009); Lerman et al. (2001) 1000
@@
Table Join semantic-join (SJ) Predict semantic join between two tables Acc He et al. (2015); Abedjan et al. (2016); Dargahi Nobari and Rafiei (2024); Wang and He (2017) 131
@@
equi-join-detect (EJ) Predict equi-joins between a set of tables F1 Lin et al. (2023); Chen et al. (2014); Rostin et al. (2009); Jiang and Naumann (2020) 517
@@
Column Transform program-transform-by-example (PTBE) Program transformation by input/output examples Acc He et al. (2018); Gulwani (2011); Devlin et al. (2017); Singh (2016) 568
@@
formula-by-context (FBC) Predict formula based on table context Acc Chen et al. (2024a, 2021b) 3626
@@
semantic-transform-by-example (STBE) Predict semantic transformations by examples Acc He et al. (2015); Abedjan et al. (2016); Dargahi Nobari and Rafiei (2024) 131
@@
Column Relationship arithmetic-relationship (AR) Predict arithmetic-relationship (AR) in tables F1 Shraga and Miller (2023); Aut 819
functional-relationship (FR) Predict functional-relationship (FR) in tables F1 Aut ; Parciak et al. (2024); Wei and Link (2019) 309
@@
string-relationship (SR) Predict string-relationship (SR) in tables F1 Aut ; He et al. (2018); Gulwani et al. (2012) 766
@@
Table understanding Needle-in-a-haystack-table (NIHT) Retrieve cell content in a table Acc Eggert et al. (2023) 1000
@@
Needle-in-a-haystack-index (NIHI) Retrieve index based on cell value Acc Eggert et al. (2023) 1000
@@
NL-2-code NL-2-SQL (NS) Translate natural-language into SQL Acc Zheng et al. (2024); Li et al. (2024); Lee et al. (2021); Yu et al. (2018); Zhong et al. (2017) 3289
@@
Table QA Table Question Answering (TQA) Answer questions based on tables Acc Wu et al. (2025); Yang et al. (2015); Chen et al. (2021c); Cheng et al. (2021) 2424
@@
Fact Verification (FV) Verify facts based on tables Acc Chen et al. (2019); Yang and Zhu (2021); Zhang et al. (2020a) 1000
@@
KB Mapping Column type annotation (CTA) Predict KB types based on column content Acc Jiménez-Ruiz et al. (2020); Yan and He (2018); Suhara et al. (2022); Zhang et al. (2019); Hulsebos et al. (2019) 1000
Column property annotation (CPA) Predict KB property for a pair of columns Acc Jiménez-Ruiz et al. (2020); Cutrona et al. (2022); Nguyen et al. (2021) 1000
Cell entity annotation (CEA) Predict KB entity for a table cell Acc Jiménez-Ruiz et al. (2020); Deng et al. (2022); Bhagavatula et al. (2015); Wu et al. (2016); Liu et al. (2023) 1000
Total 30,647
@@ 2 Related Work
Large-scale benchmarks for foundation models. The rapid advancement of foundation models has made benchmark evaluation ever more important. Prominent large-scale benchmarks, such as GLUE Wang et al. (2018) (9 NLP tasks), Super-GLUE Wang et al. (2019) (10 NLP tasks), Big-BENCH Srivastava et al. (2022) (204 tasks), MMMU Hendrycks et al. (2020) (15,908 questions), MMMU-pro Wang et al. (2024) (12,032 questions), MMLU Yue et al. (2024) (11,550 multi-modal questions) etc., offer comprehensive evaluations of model capabilities. However, as models improve rapidly, benchmarks can become saturated quickly (e.g., in the matter of a few years), prompting newer and more challenging benchmarks Wang et al. (2019, 2024). All of these benchmarking efforts have nevertheless played a crucial role in measuring and stimulating the development of foundation models.
To contribute to this growing landscape, our new MMTU benchmark comprises 30,647 challenging questions across diverse table tasks that expert users would face, which is comparable in scale with prior efforts such as MMMU and MMLU. It complements existing benchmarks, by enabling comprehensive evaluation of foundation models in the important yet underexplored area of table reasoning and understanding.
Benchmarks for reasoning. Reasoning has recently emerged as a key challenge for foundation models. Beyond general intelligence benchmarks (e.g., GPQA Diamond Rein et al. (2024), AGIEval Zhong et al. (2023), HLE ben (d)), there are specialized benchmarks targeting mathematical reasoning (e.g., AIME ben (a), MathVista ben (h), IMO ben (e)), coding (e.g., SWE-bench ben (i), CodeForce ben (c), LiveCodeBench ben (g), IOI ben (f)), and multi-modal reasoning (e.g., MMMU Yue et al. (2024), ARC ben (b)). These benchmarks have become important tools for evaluating models’ ability to tackle complex reasoning tasks, but can also get saturated quickly (e.g., GSM8k Cobbe et al. (2021), Math500 Hendrycks et al. (2021), HuamEval Chen et al. (2021a)), making it necessary to create new and more challenging benchmarks.
@@
Existing benchmarks relating to tables. Given the importance of table data, benchmarks have been developed in the ML and NLP community to evaluate model ability on tables, which however usually focus on a small set of table tasks such as NL-2-SQL Zheng et al. (2024); Li et al. (2024); Lee et al. (2021); Yu et al. (2018); Zhong et al. (2017) and Table-QA Chen et al. (2019); Wu et al. (2025); Yang et al. (2015); Chen et al. (2021c); Cheng et al. (2021). In contrast, MMTU expands the scope of current evaluations by incorporating 19 additional table tasks drawn from decades of research in communities such as data management and programming languages, resulting in a more comprehensive evaluation framework for assessing LLM capabilities on tabular data.
More recently, spreadsheet-centric benchmarks have emerged, including Spreadsheet-Bench Ma et al. (2024), SheetCopilotBench Li et al. (2023a), and Sheet-RM Chen et al. (2024b). While these are important in the spreadsheet domain, these efforts are typically limited in scale (containing a few hundred cases), and are closely tied to specific file formats (e.g., .xlsx) and software environments. In comparison, MMTU focuses on general-purpose tabular data that applies broadly across spreadsheet, database, and computational notebook settings, enabling more scalable and format-independent evaluation of foundation models.
@@ 3.1 Benchmark overview
The benchmark comprises 30,647 complex table-centric questions over 67,886 real tables, in 25 distinct task categories. Each question has a standardized format of “<Instruction, Input-Table(s), Ground-truth answer>”, like illustrated in example questions in Figure 1. Detailed statistics of the benchmark can be found in Table 2, which highlight the diversity and complexity of the questions in MMTU.
@@ 3.2 Data curation workflow
Total tables 67,886
Coding Questions 8,508 (27.8%)
 - SQL questions 3,289 (10.7%)
 - Python Pandas questions 1,593 (5.2%)
 - Spreadsheet Formula questions 3,626 (11.8%)
Non-coding Questions 22,139 (72.2%)
Questions with tables 30,647
 - Questions with 1 table 21,696 (70.8%)
 - Questions with 2 tables 5,565 (18.2%)
 - Questions with 3 or more tables 3,386 (11.0%)
@@
 - Web tables 48,502 (71.4%)
 - Spreadsheet tables 4,716 (6.9%)
 - Relational tables 14,668 (21.6%)
@@
Literature survey. To ensure our table tasks reflect real-world challenges, we draw on our experience working on related problems, that many challenging predictive table tasks have been studied in the decades worth of computer science research, particularly in data management (SIGMOD/VLDB), programming languages (PLDI/POPL), and web data (WWW/WSDM) communities. We conduct a systematic survey of publications from these venues over the past two decades, leveraging a combination of keyword searches of paper titles, and DeepResearch-like tools dee (where we specify detailed requirements for papers in these venues), to arrive at a promising set of papers and possible candidate tasks.
@@
  • (1) Real user-facing tasks, involving data tables that would otherwise require expert-level humans to perform. (We therefore exclude system-level predictive table tasks focused on performance improvements, such as query optimization Kossmann et al. (2022); Lan et al. (2021) and cardinality estimation Han et al. (2021); Kim et al. (2022));
  • (2) Objectively evaluable tasks, that come with unique manually-labeled ground truth. (We therefore exclude tasks such as table summarization Zhang et al. (2020b); Bao et al. (2018); Hancock et al. (2019); Parikh et al. (2020); Chen et al. (2013) and table augmentation Zhang and Balog (2017); Yakout et al. (2012), which lack unique ground-truth and may require subjective fuzzy LLM-based evaluations);
@@
    After the selection step, we arrive at 25 different tasks (listed in Table 1) from 52 diverse benchmark datasets (can be seen in Table 5 in the appendix). We note that a majority of these real-world tasks (all except the last 3 categories in Table 1, NL-2-code, Table-QA and KB-mapping) have not been used to evaluate foundation models. A summary of these tasks are described in more detail in Appendix A.
Data standardization and curation. To accommodate the heterogeneity across the 52 benchmark datasets (which have diverse data formats, ground-truth labels, and task definitions), we next standardize questions in each dataset into a consistent “<Instruction, Table(s), Ground-truth>” format. Figure 1 shows examples of the triplet format for different tasks. This enables consistent representation across tasks, facilitating the integration of diverse table tasks within a single benchmark framework, and allowing different LLMs to be plugged in for easy model prediction and evaluation.
Out of an abundance of caution, in this step we also prompt GPT-4o to filter out any instance of benchmark questions that may raise privacy or security concerns. Additionally, to preserve the diversity of the overall benchmark, we cap the number of questions contributed by any single dataset at 1000. The final composition of questions is reported in Table 1, with more statistics shown in Table 2 and Figure 3.
Evaluation framework. In contrast to benchmarks like MMLU Hendrycks et al. (2020) and MMMU Yue et al. (2024), which primarily use multiple-choice formats (where evaluation involves comparing a predicted option such as “A/B/C/D” against a single-letter ground truth), real-world table tasks performed by professional experts are often far more complex and nuanced. These tasks, such as code generation or structured reasoning, cannot be adequately evaluated using multiple-choice alone. In MMTU, we instead adopt a structured yet open-ended answer format (see Figure 1 for examples) for prediction and evaluation.
To support evaluations beyond simple string comparisons, we designed a lightweight evaluation framework that supports diverse evaluations in table tasks, including execution-based evaluation (for SQL and Python generation), and structured output evaluation (e.g., comparing an unordered JSON list against ground truth). Our framework is also extensible, making it easy to incorporate new task types and evaluation metrics. Details of our evaluation framework can be seen in cod .
@@
After completing all data curation steps in the workflow illustrated in Figure 2, we obtain a total of 30,647 questions that form our MMTU benchmark, with main statistics summarized in Table 2.

### v3

@@ Abstract
In this work, we introduce MMTU, a large-scale benchmark with around 28K questions across 25 real-world table tasks, designed to comprehensively evaluate models ability to understand, reason, and manipulate real tables at the expert-level. These tasks are drawn from decades’ worth of computer science research on tabular data, with a focus on complex table tasks faced by professional users. We show that MMTU require a combination of skills – including table understanding, reasoning, and coding – that remain challenging for today’s frontier models, where even frontier reasoning models like OpenAI GPT-5 and DeepSeek R1 score only around 69% and 57% respectively, suggesting significant room for improvement. We highlight key findings in our evaluation using MMTU and hope that this benchmark drives further advances in understanding and developing foundation models for structured data processing and analysis. Our code and data are available at https://github.com/MMTU-Benchmark/MMTU and https://huggingface.co/datasets/MMTU-benchmark/MMTU.
@@ 1 Introduction
Remarkable progress has been made in foundation models [foundationmodels, gpt3, r1, llm-llama, llm-palm-2], partly thanks to an expanding array of large-scale benchmarking efforts. Prominent examples include benchmarks for general language understanding (e.g., GLUE [glue], Super-GLUE [superglue], BIG-Bench [big-bench], MMLU [mmlu], MMLU-pro [mmlu-pro]), as well as benchmarks focused on coding and STEM reasoning (e.g., SWE-bench [benchmark-swebench], GPQA-diamond [benchmark-gpqa], AIME [benchmark-aime], LiveCodeBench [benchmark-livecodebench]). These efforts have played a critical role in deepening our understanding and accelerating the progress of foundation models.
Tables and table-based use cases are central to many real-world applications, including spreadsheets [excel, Google-sheets], databases [database-oracle, database-sql-server], and computational notebooks [jupyter, databricks-notebook], which often require expert-level users such as data engineers, analysts, and database administrators to operate. While LLMs have shown great promise in working with tables [excel-copilot, database-copilot, gemini-google-bigquery], existing evaluations of table-tasks remain narrow in scope – primarily focusing on NL-to-SQL [ns-bird, ns-archer, ns-spider] and Table-QA [tqa-wikiqa, tqa-hitab, tqa-tabfact] – and fail to reflect the broader spectrum of real-world tasks that professional users face.
@@
In this work, we introduce MMTU, a large-scale and challenging benchmark comprising 28,136 questions across 25 real-world table tasks. It is designed to comprehensively evaluate models’ ability to understand, reason, and manipulate real tables at the expert-level. The tasks we collect in MMTU are drawn from decades of computer science research on tabular data, with a particular focus on complex tasks that professional users face, as illustrated in Figure 1.
Our evaluation shows that the complex and technical nature of MMTU demands a combination of capabilities – including table understanding, reasoning, and coding – that remain challenging for today’s frontier models. Even the top performing models, such as OpenAI GPT-5 and DeepSeek-R1 , achieve only 69.6% and 57.9% on MMTU, suggesting substantial room for improvement, and highlighting MMTU as a strong testbed for models aspiring toward general human-level intelligence, e.g., to meet or surpass the top 10% of skilled adults in diverse technical tasks like those characterized in [agi-levels].
@@
  • Reasoning-focused models, such as OpenAI GPT-5 and DeepSeek R1, show a clear advantage over general-purpose chat models like GPT-5-Chat and DeepSeek-V3. The top reasoning models outperform the top chat models by over 10 percentage points (Table 3), underscoring the complexity of the tasks (which often require coding in SQL/Pandas) and the importance of reasoning in MMTU.
@@
Table Transform table-transform-by-relationalization (TTBR) Relationalize a table using transformation Acc [ttbr-auto-tables, transform-table-barowy2015flashrelate, transform-table-foofah, xiong2025tablecanoniser, autoprep] 230
@@
table-transform-by-output-schema (TTBS) Synthesize transformation by output schema Acc [tbps-auto-pipeline, sharma2023automatic, sharma2025datamorpher] 685
@@
table-transform-by-output-table (TTBT) Synthesize transformation by input/output tables Acc [ttbt-autopandas, ttbt-scythe, transform-table-qbo, table-transform-sql-by-output] 86
@@
Table Matching Entity matching (EM) Match rows refer to the same semantic entity Acc [em-magellandata, em-deepmatcher, em-auto-em, em-ditto, em-book] 4228
@@
Schema matching (SM) Match columns refer to the same concept F1 [sm-valentine, sm-smurf, sm-other-1, schema-matching-cupid] 669
@@
Head value matching (HVM) Match column-headers with cell-values Acc [table-gpt, schema-mapping-survey] 900
@@
Data Cleaning data-imputation (DI) Predict missing values in tables Acc [webtables, tablib, table-gpt] 1971
@@
error-detection (ED) Detect erroneous cells in tables F1 [error-detection-survey, error-detect-uni-detect, error-detect-baran, error-detect-auto-test] 1891
@@
list-to-table (L2T) Split lists of undelimited values into table Acc [l2t-tegra, list-extraction-2, list-extraction-google, list-other-1] 957
@@
Table Join semantic-join (SJ) Predict semantic join between two tables Acc [sj-sema-join, sj-dataxformer, semantic-join-dtt, auto-map] 130
@@
equi-join-detect (EJ) Predict equi-joins between a set of tables F1 [ej-auto-bi, ej-chen2014fast, ej-fk-ml, ej-hpi] 494
@@
Column Transform program-transform-by-example (PTBE) Program transformation by input/output examples Acc [ptbe-tde, ptbe-transform-prose, transform-column-devlin2017robustfill, transform-column-singh2016blinkfill, he2018transform] 558
@@
formula-by-context (FBC) Predict formula based on table context Acc [fbc-auto-formula, chen2021spreadsheetcoder] 3513
@@
semantic-transform-by-example (STBE) Predict semantic transformations by examples Acc [sj-sema-join, sj-dataxformer, semantic-join-dtt] 131
@@
Column Relationship arithmetic-relationship (AR) Predict arithmetic-relationship (AR) in tables F1 [explain-relationship, Auto-relate] 818
functional-relationship (FR) Predict functional-relationship (FR) in tables F1 [Auto-relate, fd-1, fd-2] 267
@@
string-relationship (SR) Predict string-relationship (SR) in tables F1 [Auto-relate, ptbe-tde, transform-col-flashfill] 765
@@
Table understanding Needle-in-a-haystack-table (NIHT) Retrieve cell content in a table Acc [tablib] 999
@@
Needle-in-a-haystack-index (NIHI) Retrieve index based on cell value Acc [tablib] 1000
@@
NL-2-code NL-2-SQL (NS) Translate natural-language into SQL Acc [ns-archer, ns-bird, ns-kaggledbqa, ns-spider, ns-wikisql] 2489
@@
Table QA Table Question Answering (TQA) Answer questions based on tables Acc [tqa-tablebench, tqa-wikiqa, tqa-finqa, tqa-hitab] 1793
@@
Fact Verification (FV) Verify facts based on tables Acc [tqa-tabfact, tablefact-2, tablefact-3] 916
@@
KB Mapping Column type annotation (CTA) Predict KB types based on column content Acc [semtab19, cta-autotype, cta-doduo, cta-sato, cta-sherlock] 881
Column property annotation (CPA) Predict KB property for a pair of columns Acc [semtab19, cpa-2, cpa-3] 873
Cell entity annotation (CEA) Predict KB entity for a table cell Acc [semtab19, turl, cea-1, cea-2, cea-3] 892
Total 28,136
@@ 2 Related Work
Large-scale benchmarks for foundation models. The rapid advancement of foundation models has made benchmark evaluation ever more important. Prominent large-scale benchmarks, such as GLUE [glue] (9 NLP tasks), Super-GLUE [superglue] (10 NLP tasks), Big-BENCH [big-bench] (204 tasks), MMMU [mmlu] (15,908 questions), MMMU-pro [mmlu-pro] (12,032 questions), MMLU [mmmu] (11,550 multi-modal questions) etc., offer comprehensive evaluations of model capabilities. However, as models improve rapidly, benchmarks can become saturated quickly (e.g., in the matter of a few years), prompting newer and more challenging benchmarks [superglue, mmlu-pro]. All of these benchmarking efforts have nevertheless played a crucial role in measuring and stimulating the development of foundation models.
To contribute to this growing landscape, our new MMTU benchmark comprises 28,136 challenging questions across diverse table tasks that expert users would face, which is comparable in scale with prior efforts such as MMMU and MMLU. It complements existing benchmarks, by enabling comprehensive evaluation of foundation models in the important yet underexplored area of table reasoning and understanding.
Benchmarks for reasoning. Reasoning has recently emerged as a key challenge for foundation models. Beyond general intelligence benchmarks (e.g., GPQA Diamond [benchmark-gpqa], AGIEval [benchmark-agieval], HLE [benchmark-hle]), there are specialized benchmarks targeting mathematical reasoning (e.g., AIME [benchmark-aime], MathVista [benchmark-mathvista], IMO [benchmark-imo]), coding (e.g., SWE-bench [benchmark-swebench], CodeForce [benchmark-codeforce], LiveCodeBench [benchmark-livecodebench], IOI [benchmark-ioi]), and multi-modal reasoning (e.g., MMMU [mmmu], ARC [benchmark-arc]). These benchmarks have become important tools for evaluating models’ ability to tackle complex reasoning tasks, but can also get saturated quickly (e.g., GSM8k [gsm8k], Math500 [math500], HuamEval [humaneval]), making it necessary to create new and more challenging benchmarks.
@@
Existing benchmarks relating to tables. Given the importance of table data, benchmarks have been developed in the ML and NLP community to evaluate model ability on tables, which however usually focus on a small set of table tasks such as NL-2-SQL [ns-archer, ns-bird, ns-kaggledbqa, ns-spider, ns-wikisql] and Table-QA [tqa-tabfact, tqa-tablebench, tqa-wikiqa, tqa-finqa, tqa-hitab]. In contrast, MMTU expands the scope of current evaluations by incorporating 19 additional table tasks drawn from decades of research in communities such as data management and programming languages, resulting in a more comprehensive evaluation framework for assessing LLM capabilities on tabular data.
More recently, spreadsheet-centric benchmarks have emerged, including Spreadsheet-Bench [ma2024spreadsheetbench], SheetCopilotBench [li2023sheetcopilot], and Sheet-RM [chen2024sheetagent]. While these are important in the spreadsheet domain, these efforts are typically limited in scale (containing a few hundred cases), and are closely tied to specific file formats (e.g., .xlsx) and software environments. In comparison, MMTU focuses on general-purpose tabular data that applies broadly across spreadsheet, database, and computational notebook settings, enabling more scalable and format-independent evaluation of foundation models.
@@ 3.1 Benchmark overview
The benchmark comprises 28,136 complex table-centric questions over 61,763 real tables, in 25 distinct task categories. Each question has a standardized format of “<Instruction, Input-Table(s), Ground-truth answer>”, like illustrated in example questions in Figure 1. Detailed statistics of the benchmark can be found in Table 2, which highlight the diversity and complexity of the questions in MMTU.
@@ 3.2 Data curation workflow
Total tables 61,763
Coding Questions 7,331 (26.1%)
  - SQL questions 2,489 (18.8%)
  - Python Pandas questions 1,329 (4.5%)
  - Spreadsheet Formula questions 3,513 (11.9%)
Non-coding Questions 20,805 (73.9%)
Questions with tables 28,136
  - Questions with 1 table 20,048 (71.3%)
  - Questions with 2 tables 5,285 (18.8%)
  - Questions with 3 or more tables 2,803 (10.0%)
@@
  - Web tables 46,264 (74.9%)
  - Spreadsheet tables 4,540 (7.4%)
  - Relational tables 10,959 (17.7%)
@@
Literature survey. To ensure our table tasks reflect real-world challenges, we draw on our experience working on related problems, that many challenging predictive table tasks have been studied in the decades worth of computer science research, particularly in data management (SIGMOD/VLDB), programming languages (PLDI/POPL), and web data (WWW/WSDM) communities. We conduct a systematic survey of publications from these venues over the past two decades, leveraging a combination of keyword searches of paper titles, and DeepResearch-like tools [deepresearch] (where we specify detailed requirements for papers in these venues), to arrive at a promising set of papers and possible candidate tasks.
@@
  • (1) Real user-facing tasks, involving data tables that would otherwise require expert-level humans to perform. (We therefore exclude system-level predictive table tasks focused on performance improvements, such as query optimization [qo-1, qo-2] and cardinality estimation [cardinality-1, cardinality-2]);
  • (2) Objectively evaluable tasks, that come with unique manually-labeled ground truth. (We therefore exclude tasks such as table summarization [table-summary-2, table-summary-3, table-summary-google, table-summary-4, table-summary-5] and table augmentation [table-augmentation-2, table-augmentation-infogather], which lack unique ground-truth and may require subjective fuzzy LLM-based evaluations);
@@
    After the selection step, we arrive at 25 different tasks (listed in Table 1) from 52 diverse benchmark datasets (can be seen in Table 5 in the Appendix). We note that a majority of these real-world tasks (all except the last 3 categories in Table 1, NL-2-code, Table-QA and KB-mapping) have not been used to evaluate foundation models. A summary of these tasks is described in more detail in Appendix B.
Data standardization and curation. To accommodate the heterogeneity across the 52 benchmark datasets (which have diverse data formats, ground-truth labels, and task definitions), we next standardize the questions in each dataset into a consistent “<Instruction, Table(s), Ground-truth>” format. Figure 1 shows examples of the triplet format for different tasks. This enables consistent representation across tasks, facilitating the integration of diverse table tasks within a single benchmark framework, and allowing different LLMs to be plugged in for easy model prediction and evaluation.
To ensure the quality of MMTU, we conduct an LLM-based quality check using o4-mini. The model is prompted to evaluate each question for two aspects: (1) whether the question is ambiguous (since tasks like NL-code can sometimes be inherently ambiguous [saparina2024ambrosia]), and (2) whether the ground-truth answer is correct. Approximately 8% of the original questions were flagged as either ambiguous or incorrect and subsequently removed to enhance benchmark quality.
As an additional safeguard, we also use LLM to exclude any benchmark instances that might pose privacy or security risks. To maintain diversity across sources, we cap the number of questions drawn from any single dataset at 1000.
The final composition of the questions is reported in Table 1, with more statistics shown in Table 2 and Figure 3.
Evaluation framework. In contrast to benchmarks like MMLU [mmlu] and MMMU [mmmu], which primarily use multiple-choice formats (where evaluation involves comparing a predicted option such as “A/B/C/D” against a single-letter ground truth), real-world table tasks performed by professional experts are often more complex and nuanced. These tasks, such as code generation or structured reasoning, cannot be adequately evaluated using multiple-choice alone. In MMTU, we instead adopt a structured yet open-ended answer format (see Figure 1 for examples) for prediction and evaluation.
To support evaluations beyond simple string comparisons, we designed a lightweight evaluation framework that supports diverse evaluations in table tasks, including execution-based evaluation (for SQL and Python generation), and structured output evaluation (e.g., comparing an unordered JSON list against ground truth). Our framework is also extensible, making it easy to incorporate new task types and evaluation metrics. Details of our evaluation framework can be seen in [code-data].
@@
After completing all data curation steps in the workflow illustrated in Figure 2, we obtain a total of 28,136 questions that form our MMTU benchmark, with main statistics summarized in Table 2.